Solve each of the following quadratic equations by completing the square.

2x^2+10x+11=0

Answer by checkley75(3666)

2x^2+10x+11=0
using the quadratic equation we get
X=(-10+-SQRT[10^2-4*2+11])/2*2
X=(-10+-SQRT[100-88])/4
X=(-10+-SQRT12)/4
X=(-10+-3.464)/4
X=(-10+3.464)/4
X=-6.536/4
X=-1.634 ANSWER.
X=(-10-3.464)/4
X=-13.464/4
X=-3.366 ANSWER.